Solve the given equations by factoring.

5a. x^2-14=5x

5b. 6x^2+7x+2=0
1
Expert's answer
2013-01-21T11:17:08-0500
5a. x^2-14=5xx^2-5x-14=0
x^2-7x+2x-14=0
x(x-7)+2(x-7)=0
(x+2)(x-7)=0
so x1=-2, x2=7
5b. 6x^2+7x+2=06x^2+3x+4x+2=0
3x(2x+1)+2(2x+1)=0
(3x+2)(2x+1)=0
x=-2/3, x=-1/2